OpenStudy (anonymous):

Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 4)(x − 4) and (x + 4)(x + 4)

OpenStudy (anonymous):

@Mertsj @jim_thompson5910

OpenStudy (anonymous):

@DollyAcquah

OpenStudy (anonymous):

Someone, help!

OpenStudy (mertsj):

The two factors are the same so the answer to the two problems will be the same. That is because multiplication is a commutative operation which means that the order of the factors does not matter. The answer will be the same.

OpenStudy (anonymous):

Thank you so much! @Mertsj

OpenStudy (mertsj):

yw

OpenStudy (nikato):

Wait, no. For (x+4)(x-4)= x^2 -16 (x+4)(x+4)= x^2 + 8x +16 They are two different answers They both have x^2 However the first is -16 but the second has +8x+16

OpenStudy (nikato):

The first one also doesn't have "a middle term"

OpenStudy (mertsj):

Oh gee. I didn't see that plus sign. thought it was a minus sign.

OpenStudy (mertsj):

The first one is the product of two binomials whose outer and inner products are opposites. Therefore the result is a binomial. The second one is the product of two identical factors and the result is a trinomial square.